OpenStudy (anonymous):

What are the roots for the function x^2+4x-3=0

OpenStudy (campbell_st):

well you need the general quadratic formula of you can complete the square... which do you find easier...?

OpenStudy (anonymous):

the answers given was a. x= -2,6 b. x= -4, 12 c. x= -2+-sqrt(7) d. x=2+-21sqrt(2) I'm sure a and b is out. How about c and d? how do i do this when there is sqrt?

OpenStudy (campbell_st):

well its quite simple... and no need to guess... you can solve by 1. completing the square 2. using the general quadratic formula it depends on which you are more comfortable with ..

OpenStudy (anonymous):

note, please don't get confused, +- means the plus is on top and the minus is below the + sign

OpenStudy (anonymous):

i prefer the completing the square method

OpenStudy (campbell_st):

ok... so rewrite the quadratic as \[x^2 + 4x = 3\] what do you need to add to the left hand side to get a perfect square..?

OpenStudy (anonymous):

im not sure... A bit confusing

OpenStudy (campbell_st):

ok... so a perfect square is \[(x +b)^2 = x^2 + 2bx + b^2\] so what you need to do is halve the coefficient of the middle term then square it.. so in you question what is half of 4 squared...?

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

well have of the coefficient is 2 then squared is 4 so you need to add 4 to both sides of the equation \[x^2 + 4x + 4 = 3 + 4\] this can be simplified to \[(x + 2)^2 = 7\] so now all you need to do is solve for x. hope this helps.

OpenStudy (anonymous):

what if i want to use the the general quadratic formula?

OpenStudy (campbell_st):

you'll get the same result you can also get a clue using the discriminant... if you know what it is

OpenStudy (campbell_st):

so do you know how to solve it from here...?

OpenStudy (anonymous):

yes i just solved it using the completing the square which is the answer is c

OpenStudy (anonymous):

i tried using the general quadratic formula but i didt get the answer

OpenStudy (campbell_st):

ok... so lets get back to completing the square \[(x + 2)^2 = 7\] whats the 1st step

OpenStudy (campbell_st):

and your solution using the GQF maybe caused by not knowing how to simplify a radical

OpenStudy (anonymous):

i would like to know when do u use this formula \[(x+b)^{2}=x ^{2}+2bx+b ^{2}\]? because i only recall how to do (x+_) (x-_). When must this be used and when must that formula u given me be used?

OpenStudy (campbell_st):

the left hand side is a perfect square, the right hand side is the expansion. the left hand side is (x + b)(x + b)

OpenStudy (campbell_st):

what you recall is the difference of 2 squares \[(x -a)(x + a) = x^2 - a^2\]

OpenStudy (campbell_st):

ok, so lefts solve the equation above... whats the 1st step

OpenStudy (campbell_st):

its \[(x + 2)^2 = 7\]

OpenStudy (campbell_st):

any idea on how to start solving the equation...?

OpenStudy (anonymous):

yes i solved it. again the general quadratic formula i didt get it. the answer is not the same

OpenStudy (campbell_st):

all you need to do with the complete the square method is \[(x +2)^2 = 7\] take the square root of both sides \[x + 2 = \pm \sqrt{7}\] subtract 2 from both sides \[x = -2 \pm \sqrt{7}\]

OpenStudy (anonymous):

i got it... thank u campbell... you were so helpful